OpenStudy (anonymous):

Express 12+√-81 in terms of i. A.21+i B.21i C.12-9i D.12+9i

OpenStudy (anonymous):

\[\sqrt{-81}=\sqrt{81}\times \sqrt{-1}=9i\]

OpenStudy (unklerhaukus):

D.

OpenStudy (anonymous):

so last one

OpenStudy (anonymous):

SATELLITE! I need ur help!